Chandria:

the size of basketball floors varies due to building sizes and other concerns such as cost etc. the length L is to be at most 94ft and the width W is to be at most 50ft a. write a system of linear inequalities that describes the situation in tye problem b. graph the system that describes the possible dimensions of a basketball floor

Chandria:

Please help me 😥

Chandria:

Write four systems of four inequalities that describe a 3-unit-by-3-unit square that has (0,0) as one of the vertices.

AZ:

So for your original question, when we say length is "at most" 94 think about it, it can be AT MOST 94 so it can be 94 or it can be anything less than 94 would that be < or \(\le\) L < 94 or L \(\le\) 94 < is less than \(\le\) is less than or equal to

AZ:

similarly So for your original question, when we say width is "at most" 50 feet think about it, it can be AT MOST 50 so it can be 94 or it can be anything less than 50 would that be < or \(\le\) W < 50 or W \(\le\) 50 < is less than \(\le\) is less than or equal to

AZ:

you can use something like https://www.desmos.com/calculator to graph it

Chandria:

How can I graph it ?

AZ:

You need to first determine if it's \( <\) or \(\le\) so which one do you think it is? remember, we're saying AT MOST so that means it does include that number and everything below it is that going to be \( <\) (less than) or would that be the \(\le\) (less than or equal to) sign?

AZ:

To graph it, you can use Desmos but instead of using the letters 'L' and 'W' just use 'x' because then you'll be able to see the graph